Given πœ‡=72; and 𝜎=8, find the z-score that corresponds to each of the following scores up to two decimal place


1
Expert's answer
2022-03-16T16:09:52-0400

a.


"z=\dfrac{x-\mu}{\sigma}=\dfrac{68-72}{8}=-0.50"

b.


"z=\dfrac{x-\mu}{\sigma}=\dfrac{80-72}{8}=1.00"